# The effect of diversity on group decision-making

### v1

@@ Abstract
Code and data used for the analysis are available in the anonymised repository: https://anonymous.4open.science/r/cogsci24-FD6D.
@@ Introduction
Group deliberation is the dialogue between a group that facilitates the careful weighing of information, providing adequate speaking opportunities, and bridging the differences among participants’ diverse ways of speaking and knowing (?, ?). ? (?) review evidence that, in the right conditions, groups can outperform even the most successful individual members in their decision-making. These conditions include common goals among group members and the exchange of arguments. As well as groups exceeding the capacity of any individual member, another relevant standard is the decision which would be reached by simple aggregation or averaging of independent individual judgements (the so-called “Wisdom of the Crowd” effect, ?, ?). In the study of ? (?) small groups outperform the wisdom of the crowd standard. ? (?) argue that small groups can be more effective than both individuals of superior ability, and large crowds. Further, this and other evidence (?, ?, ?) supports the idea that the exchange of arguments is mechanistically related to successful group decision-making.
That said, group decision-making is not universally successful. A tradition in social psychology has emphasised the potential for negative outcomes. A notable diagnosis is the perils of “Group Think” (?, ?). Potential mechanisms of failures of group decision-making include social loafing, herding, deference to others’ preferences — real (?, ?), or imagined (?, ?) —, failure to share knowledge available to individual group members (?, ?), underweighting dissenting voices, and amplification of individual cognitive biases.
We note that the mechanisms supporting the Wisdom of Crowds effect relies on individuals having independent access to some true information (signal), and uncorrelated errors (noise). To the extent that individuals share systematic error, such as a shared cognitive bias, aggregation of individual judgements will reinforce error, not eliminate it.
@@
In this work, we leverage a unique dataset of group decisions, the DeliData corpus (?, ?). In DeliData, 500 groups provided initial judgments and then discussed the Wason Card Selection Task, for which full transcripts were recorded. Individuals then submitted their updated answers. These data allow us to look, with a high degree of granularity, at the features of groups and the deliberative processes which generated successful and unsuccessful deliberation. Furthermore, DeliData contains data indicating participants’ solutions, arguments, and deliberative markers, which would be beneficial for rich and multifaceted analysis.
@@
Second, we hypothesise that a greater diversity of initial ideas will support better group performance. Following ? (?) we index this by inspecting the number of distinct individual solutions submitted by participants prior to entering the group discussion. Third and finally, that diversity in discussion content will support better group performance. This was shown for forecasting by ? (?). We index this based on the distinct solutions mentioned in the conversation, as well as the transitions between mentioned solutions.
@@
To create the dataset, the Wason Card Selection Task was presented in its traditional form: it is explained that cards have a number on one side and a letter on the other. Four cards are shown displaying an even number (‘E’), odd number (‘O’), consonant (‘C’) and vowel (‘V’). The task is to answer the question ”Which card(s) should you turn to test the rule: All cards with vowels on one side, have an even number on the other”. The intuitive but wrong answer (henceforth the Lure) is to turn the even number and vowel (‘EV’). The correct answer (henceforth the Target) is to turn over the odd number and the vowel (‘OV’).
@@
The Wason Card Selection task has one correct answer, allowing for a direct and objective measure of performance. In this work, we leverage the fine-grained scoring from (?, ?), where 0.25 points each are given for i) selecting either card in the Target (O and V) and (ii) for not turning unnecessary cards (E and C). Therefore, if the participant submitted a correct solution, their score would be 1, if they are off by one card - 0.75 and so on. If participants selected the Lure (EV), their fine-grained score would be 0.5. Similarly, selecting all cards (CEOV) or none, will have a fine-grained score of 0.5.
@@
Finally, we analyse the interaction of the measures above and the probing utterances, as annotated in DeliData (?, ?)
@@ Results
As a first step of this analysis, we are hypothesising that group size is a factor that contributes towards better cognitive diversity and ultimately to improved decision-making performance (as measured by performance gain).
Considering only the size of the group and the performance gain, we show that the performance gain of a group increases consistently with group size, as shown in Figure 1. To confirm the significance of these results, we perform a 1-sample t-test against the baseline performance gain of 0. This yields p-values of 2.5-8, 3.5-15, 4.1-12, and 2.8-10 for groups of 2, 3, 4, 5 respectively. This indicates that for all group sizes, the performance gain was significantly above 0. Following, we performed an ANOVA (?, ?) comparison between groups of sizes 2 to 5. This showed a significant result (p=0.0251), confirming that group size affects performance gain. Following this, we did a post-hoc analysis, performing a t-test between different group sizes. The results (also presented in table 2) of the test show that there is no significant difference between each of the groups. This is also consistent and confirms the findings of (?, ?). That said, ? (?) argues that groups (3+ people) perform better than dyads (dialogues). To this end, we also confirm this claim on the 500 dialogues of DeliData, showing that groups (size>=>=> =3) significantly outperform dyads (size=2) with a p-value of 0.03.
@@
Many factors can contribute towards the difference between dyads (2 participants) and multi-party dialogues (3 or more participants). First, larger groups may benefit from an increased idea pool, which we investigate in detail in the next section - Diversity of initial ideas. Secondly, larger groups may engage in more diverse conversational strategies, thus incorporating more ideas and arguments during the group discussion. Figure 2 compares basic conversational statistics between dyads and groups. Dialogues between two interlocutors have mostly between 10 and 25 utterances, while group dialogues have a wider range of lengths, with a long tail of dialogues longer than 50 utterances. Likewise, groups in these dialogues tend to use a larger vocabulary, as shown on the violin plot of the unique tokens. We further evaluate conversational patterns and characteristics in section Diversity of discussed solutions.
@@
This analysis shows that most of the predictive power comes from the diversity of initial solutions, and including the number of participants does not bring additional predictive power. Another perspective would be that having a higher diversity of initial solutions already captures the information of having a larger group size.
@@
Pearson Correlation p-value
dialogue length 0.10 0.02
unique solutions 0.28 2.8-10
unique transitions 0.22 4.01-7
stuck transitions 0.02 0.67
circular transitions 0.19 8.99-6
Table 3: Results of correlation analysis between different factors representing the diversity of solutions and performance gain. In bold, we show the statistically significant results.
Here we use the five indicators outlined in the Analytic strategy subsection - dialogue length, number of unique solutions, number of unique transitions, stuck transitions and circular transitions. In order to evaluate these five indicators and their relationship to performance gain, we perform a correlation study using (?, ?). The values of the coefficients and the statistical significance values are presented in Table 3. Out of the five factors, four are statistically significant in predicting the performance gain - the dialogue length, the number of unique discussed solutions, the number of unique transitions, and the number of circular transitions. Further, if we consider the correlation coefficients, the number of unique solutions (0.28) is considered the most important feature, followed by the number of unique transitions (0.22), and the circular transitions (0.19).
@@
In the previous subsection (Diversity of initial ideas), we showed that groups with a high diversity of initial solutions achieve better performance gain. Likewise, in this section, we showed that the number of discussed solutions is also associated with increased performance gain. However, not all groups start with a diverse set of initial ideas. A reasonable research question would be - what happens in conversations where groups with low initial diversity achieve a high number of discussed ideas. To evaluate this, we test the hypothesis that conversational probing affects dialogues with a low diversity of initial ideas. The premise is that if participants probe the conversation more often, this will push the group to discuss novel ideas. Here, we investigate the correlation between the number of conversational probes and the diversity of discussed solutions. The results of this analysis are presented in figure 4, and show that there is a weak correlation between the number of probing utterances and the increase of unique solutions discuss.
Figure 4: Scatter plot of relation between unique solutions discussed and number of probing utterances, for conversations with low diversity of initial ideas.
To further analyse this effect, we take advantage of the more granular probing annotations in DeliData, namely whether the probing was related to Moderation, Reasoning behind a solution, or whether a specific Solution should be picked. To that end, we perform a correlation study to investigate the relationship between different kinds of probing and the diversity of discussed solutions. Results of this analysis are shown in Table 4. We see that both probing for reasons and solutions are a significant contributor towards predicting the diversity of discussed solutions (p-values $\approx$ 0). The correlation coefficients show that probing for reasons (0.41) has a higher association with the diversity of discussed solutions than simply enquiring about solutions (PearsonR = 0.26). This is an interesting result, showing the importance of reasoning and also confirming the findings of (?, ?) about the value of deliberation. Moreover, this gives a basis for future work on interventional analysis and applications, where conversational probing can increase cognitive diversity and improve conversational outcomes.
Pearson Correlation p-value
Moderation -0.017 0.71
Reasoning 0.41 3.24-21
Solution 0.26 4.21-9
Table 4: Results of correlation analysis of different types of probing and predicting the number of unique solutions discussed. In bold, we show the statistically significant results.
@@ Discussion
On the chosen task, which has a high risk of individual bias and error, online discussion groups outperform individuals, and further groups of 3 and above outperform dyads.
@@
This analysis of the DeliData corpus shows that it replicates the finding that group discussion of the Wason Task can reduce, rather than reinforce, the individual susceptibility to error on the task. It extends this finding to the increasingly relevant domain of online discussion among strangers. The finding is all the more remarkable for the low-commitment  [^2] As in DeliData the participants were recruited from the crowd-sourcing website MechanicalTurk, brief, nature of the discussions (typically only 20 - 40 utterances).

### v2

@@ Abstract
Code and data used for the analysis are available in the repository: https://github.com/gkaradzhov/cognitive-diversity-groups-cogsci24.
@@ Introduction
Group deliberation is the dialogue that facilitates the careful weighing of information, providing adequate speaking opportunities, and bridging the differences among participants’ diverse ways of speaking and knowing (?, ?). ? (?) review evidence that, in the right conditions, groups can outperform even the most capable individual members in their decision-making. These conditions include common goals among group members and the exchange of arguments. As well as groups exceeding the capacity of any individual member, another relevant standard is the decision which would be reached by simple aggregation or averaging of independent individual judgements (the so-called “Wisdom of the Crowd” effect, ?, ?). In the study of ? (?) small groups outperform the wisdom of the crowd standard. ? (?) argue that small groups can be more effective than both individuals of superior ability, and large crowds. Further, this and other evidence (?, ?, ?) supports the idea that the exchange of arguments is essential to successful group decision-making.
That said, group decision-making is not universally successful. A tradition in social psychology has emphasised the potential for negative outcomes. A notable diagnosis is the perils of “Group Think” (?, ?). Potential mechanisms of failures of group decision-making include social loafing, herding, deference to others’ preferences — real (?, ?), or imagined (?, ?) —, failure to share knowledge available to individual group members (?, ?), underweighting dissenting voices, and amplification of individual cognitive biases. We note that the mechanisms supporting the Wisdom of Crowds effect relies on individuals having independent access to some true information (signal), and uncorrelated errors (noise). To the extent that individuals share systematic error, such as a shared cognitive bias, aggregation of individual judgements will reinforce error, not eliminate it.
@@
In this work, we leverage a unique dataset of group decisions, the DeliData corpus (?, ?). In DeliData, 500 groups provided initial judgments and then discussed the Wason Card Selection Task, for which full transcripts were recorded. Individuals then submitted their updated answers. This dataset allows us to look, with a high degree of granularity, at the features of groups and the deliberative processes which generated successful and unsuccessful deliberation. Furthermore, DeliData contains data indicating participants’ solutions, arguments, and deliberative markers, which would be beneficial for rich and multifaceted analysis.
@@
Second, we hypothesise that a greater diversity of initial ideas will support better group performance. Following ? (?) we index this by inspecting the number of distinct individual solutions submitted by participants prior to entering the group discussion. Third and finally, that diversity in discussion content will support better group performance. This was shown for forecasting by ? (?). We index this based on the distinct solutions mentioned in the conversation, as well as the transitions between mentioned solutions. We further hypothesise, that conversational probing (i.e. asking questions that provoke further discussion (?, ?)) has an effect over the diversity of discussion content. The premise is that conversational probing can be considered as a possible intervention that would increase the content and cognitive diversity of a discussion.
@@
To create the dataset, the Wason Card Selection Task was presented in its traditional form: it is explained that cards have a number on one side and a letter on the other. Four cards are shown displaying an even number (‘E’), odd number (‘O’), consonant (‘C’) and vowel (‘V’). The task is to answer the question “Which card(s) should you turn to test the rule: All cards with vowels on one side, have an even number on the other”. The intuitive but wrong answer (henceforth the Lure) is to turn the even number and vowel (‘EV’). The correct answer (henceforth the Target) is to turn over the odd number and the vowel (‘OV’).
@@
The Wason Card Selection task has one correct answer, allowing for a direct and objective measure of performance. In this work, we leverage the fine-grained scoring from ? (?), where 0.25 points each are given for i) selecting either card in the Target (O and V) and (ii) for not turning unnecessary cards (E and C). Therefore, if the participant submitted a correct solution, their score would be 1, if they are off by one card - 0.75 and so on. If participants selected the Lure (EV), their fine-grained score would be 0.5. Similarly, selecting all cards (CEOV) or none, will have a fine-grained score of 0.5.
@@
We conclude this section by also analysing the interaction of the measures above and the probing utterances, as annotated in DeliData (?, ?)
@@ Results
As a first step of this analysis, we are hypothesising that group size is a factor contributing to better cognitive diversity and ultimately improved decision-making performance (as measured by performance gain).
Considering only the size of the group and the performance gain, we show that the performance gain of a group increases consistently with group size, as shown in Figure 1. To confirm the significance of these results, we perform a 1-sample t-test against the baseline performance gain of 0. This yields p-values of 2.5-8, 3.5-15, 4.1-12, and 2.8-10 for groups of 2, 3, 4, 5 respectively. This indicates that for all group sizes, the performance gain was significantly above 0. Next, we performed an ANOVA (?, ?) comparison between groups of sizes 2 to 5. This showed a significant result (p=0.0251), confirming that group size affects performance gain. Following this, we did a post-hoc analysis, performing a t-test between different group sizes. The results (also presented in Table 2) of the test show that there is no significant difference between each of the groups. This is also consistent and confirms the findings of (?, ?). That said, ? (?) argues that groups (3+ people) perform better than dyads (dialogues). To this end, we also confirm this claim on the 500 dialogues of DeliData, showing that groups (size>=>=> =3) significantly outperform dyads (size=2) with a p-value of 0.03.
@@
Overall, we find that group size has an effect on performance gain. However, group size being a proxy measure, we cannot directly test whether the increase in performance gain is due to increased cognitive diversity or other factors. That said, in this work, we further investigate two characteristics that measure cognitive diversity directly. First, larger groups may benefit from an increased idea pool, which we investigate in detail in the next section - Diversity of initial ideas. Secondly, larger groups may engage in more diverse conversational strategies, thus incorporating more ideas and arguments during the group discussion. Figure 2 compares basic conversational statistics between dyads and groups. Dialogues between two interlocutors have mostly between 10 and 25 utterances, while group dialogues have a wider range of lengths, with a long tail of dialogues longer than 50 utterances. Likewise, groups in these dialogues tend to use a larger vocabulary, as shown on the violin plot of the unique tokens. We further evaluate conversational patterns and characteristics in section Diversity of discussed solutions.
@@
This analysis shows that most of the predictive power comes from the diversity of initial solutions, and including the number of participants does not bring additional predictive power. Notably, another perspective would be that having a higher diversity of initial solutions already captures the information of having a larger group size.
@@
Correlation PCR coef.
dialogue length 0.10 0.01
unique solutions 0.28 0.02
unique transitions 0.22 0.02
stuck transitions 0.02 -0.02
circular transitions 0.19 0.01
Table 3: Results of correlation and PCR analyses between different factors representing the diversity of solutions and performance gain. In bold, we show the statistically significant values with p-value <<< 0.05.
Here we use the five indicators outlined in the Analytic strategy subsection - dialogue length, number of unique solutions, number of unique transitions, stuck transitions and circular transitions. In order to evaluate these five indicators and their relationship to performance gain, we perform (i) a correlation study using (?, ?) and (ii) a regression analysis using Principal Component Regression (PCR) (?, ?). For the latter, we transform the input feature space using principal component analysis. This can handle adequately the multicollinearity present among the features in the dataset. We implement it by fitting a linear regression over the transformed feature space. We then transform the coefficients of the regression back to the original feature space to obtain the final coefficients.
The values of the correlation analysis and the PCR analysis are presented in Table 3. If we consider the correlation statistic, out of the five factors, four are statistically significant in predicting the performance gain - the dialogue length, the number of unique discussed solutions, the number of unique transitions, and the number of circular transitions. Further, if we consider the correlation coefficients, the number of unique solutions (0.28) is considered the most important feature, followed by the number of unique transitions (0.22), and the circular transitions (0.19). Similarly, if we analyse the results of the PCR analysis, dialogue length, number of unique solutions, and circular transitions are significant. Similarly to the correlation analysis, the number of unique solutions has the highest importance, based on the coefficients from the principal component regression.
@@
In the previous subsection (Diversity of initial ideas), we showed that groups with a high diversity of initial solutions achieve better performance gain. Likewise, in this section, we showed that the number of discussed solutions is also associated with increased performance gain. However, not all groups start with a diverse set of initial ideas. A reasonable research question would be - what happens in conversations where groups with low initial diversity achieve a high number of discussed ideas. To evaluate this, we test the hypothesis that conversational probing affects dialogues with a low diversity of initial ideas. The premise is that if participants probe the conversation more often, this will push the group to discuss novel ideas. To this end, we calculate a Pearson-R correlation of 0.34 between the number of conversational probes and the diversity of discussed solutions.
To further analyse this effect, we take advantage of the more granular probing annotations in DeliData, namely whether the probing was related to Moderation, Reasoning behind a solution, or whether a specific Solution should be picked. To that end, we perform correlation and PCR studies to investigate the relationship between different kinds of probing and the diversity of discussed solutions. Results of this analysis are shown in Table 4. We see that both probing for reasons and solutions are a significant contributor towards predicting the diversity of discussed solutions. The correlation coefficients show that probing for reasons (0.41) has a higher association with the diversity of discussed solutions than simply enquiring about solutions (PearsonR = 0.26). Likewise, the PCR coefficient for probing for reasons is 0.83, and probing for solutions is 0.43. This indicates that regardless of the type of analysis (correlational and predictive), we show the importance of probing for reasons, followed by probing for solutions. This is an interesting result, showing the importance of reasoning and also confirming the findings of (?, ?) about the value of deliberation. Moreover, this gives a basis for future work on interventional analysis and applications, where conversational probing can increase cognitive diversity and improve conversational outcomes.
Correlation PCR Coef.
Moderation -0.017 -0.06
Reasoning 0.41 0.83
Solution 0.26 0.43
Table 4: Results of correlation and PCR analyses of different types of probing and predicting the number of unique solutions discussed. In bold, we show the statistically significant values.
@@ Discussion
On the chosen task, which has a high risk of individual bias and error, online discussion groups outperform individuals, and groups of 3 and above outperform dyads.
@@
This analysis of the DeliData corpus shows that it replicates the finding that group discussion of the Wason Task can reduce, rather than reinforce, the individual susceptibility to error on the task. It extends this finding to the increasingly relevant domain of online discussion among strangers. The finding is all the more remarkable for the low-commitment [^2] In DeliData the participants were recruited from the crowd-sourcing website MechanicalTurk and the brief nature of the discussions (typically only 20 - 40 utterances).
@@ Conclusion
Acknowledgements
Georgi Karadzhov’s PhD was supported by an EPSRC doctoral training scholarship. This research was additionally supported by a donation from Google.